OpenStudy (anonymous):

At 700 K, CCl4 decomposes to carbon and chlorine. The Kp for the decomposition is 0.76. Find the starting pressure of CCl4 at this temperature that will produce a atotal pressure of 1.9atm at equilibrium.

OpenStudy (surry99):

Start with a balanced chemical equation and then set up your ICE table

OpenStudy (cuanchi):

This is not an easy problem you have to do a system of 2 equations with 2 unknowns 1) is the Kp = 0.76 the other is the total pressure 2) Pt= 1.9 atm = Pccl4 + Pc + Pcl2 if you do the ICE table you will find the pressures at the equilibrium in function of the initial pressure CCl4 -> C + 2 Cl2 Kp= Pc (Pcl2)^2 /Pccl4 Pccl4 = P-x Pc= x Pcl2= 2x you can replace these values in the equation 2 and isolate P 3) P= 1.9/ (2x) replace P in the equilibrium formula for 1.9/(2x) and solve for x when you get the value of x, then go back to equation 3 to calculate initial P I got this values Initial P for CCl4=1.6284 atm P of C at equilibrium = 0.5834 P of Cl at equilibrium = 1.1668 P of CCl4 at equilibrium = 1.045